6. >>x=[1 2 4 6 12]; >m mean(x) > var(x)
Answer #1

Python Program using numpy:-

---------------------------------------------------------------

import numpy
x=[1,2,4,6,12]
m=numpy.mean(x)
v=numpy.var(x)
print "Mean of the list is ",m
print "Variance of the list is ",v

--------------------------------------------------------------------

Program Output:-

Execute Embed main.py STDIN 1 import numpy 2 x=[1,2,4,6,12] 3 m-numpy.mean(x) I.lı Result $python main.py Mean of the list is

----------------------------------------------------------

Finally mean is-------5

Variance is---------15.2

--------------------------------------------------

Explanation:-

------------------------------------------

1.we import numpy to calculate Mean and Variance in python

2.Mean is simply "Average of all numbers"--------------------(1+2+4+6+12)/5=5

3.Formula for variance is below

に!

Steps:-

X1=1,

X2=2,

X3=4,

X4=6,

X5=12

------------------------

Mean(u)=5

-----------------

Number of items in the list(n)=5

----------------

Square of (1-5)=16

Square of (2-5)=9

Square of (4-5)=1

Square of (6-5)=1

Square of (12-5)=49

--------------------------------------------

Add all the elements=16+9+1+1+49=76

So finally variance=76/5=15.2

------------------------------------------------------------------------------

If you have any doubts regarding anything please comment defintly i will respond.Please please rate if you are satisfied.Thank you.